# AI for Just Work: Constructing Diverse Imaginations of AI beyond “Replacing Humans”

### v1

@@ 1. The phenomenon of monoculture in the AI community, and why questioning the objectives of AI matters?
Historical studies show that the AI field did not develop in a vacuum, but was closely related to the social, economic, and political power, ideologies, and values at a given time. Scholars have traced the linkage between AI and ideologies such as eugenics, racism, colonialism, anti-democratic and anti-humanist utopianism (Burrell & Metcalf, 2024; Ali et al., 2023). It should be noted that although the majority in the AI community may not be aware of or do not agree with these ideologies, it is the power and historical contingency that enabled a few privileged people, not the public or the majority in the community via a democratic process, to influence the research agenda and the underlying ideologies that set the paths of least resistance in the AI community. As the previous part indicates, once the rules of the game are set, we can know where the game is going without any information about the game players. And the rules of the game can appear to be necessary, obvious, natural, and inevitable. As Jenna Burrell and Jacob Metcalf put it in their editorial Introduction for the special issue of “Ideologies of AI and the consolidation of power”: Naming power (Burrell & Metcalf, 2024):
@@
A number of works have identify the goal of AI and its roots in eugenics. In their book How Data Happened: A History from the Age of Reason to the Age of Algorithms (Wiggins & Jones, 2023), Chris Wiggins and Matthew L. Jones trace the historical genesis of the contemporary statistics and data science in figures such as Francis Galton, who is the originator of eugenics, and Karl Pearson, who institutionalized eugenics. In her book Discriminating Data: Correlation, Neighborhoods, and the New Politics of Recognition (Chun, 2021), Wendy Hui Kyong Chun reveals the ties between correlation in machine learning and its eugenic history: both “seek to tie the past to the future—correlation to prediction—through supposedly eternal, unchanging biological attributes.” In their paper The TESCREAL bundle: Eugenics and the promise of utopia through artificial general intelligence (Gebru & Torres, 2024), Timnit Gebru and Émile P. Torres trace the field’s current unquestioned motivation for artificial general intelligence (AGI) back to the normative framework of the newer version of eugenics pushed by a few powerful leading figures. The eugenic ideals are dubbed the “TESCREAL bundle,” which denotes “transhumanism, Extropianism, singularitarianism, (modern) cosmism, Rationalism, Effective Altruism, and longtermism.” These utopia goals justify the creation of unsafe algorithms, concentration of power, evasion of accountability, and ongoing harms to marginalized groups upon which AI is built. Furthermore, such ideologies are naturalized so that they appear as inevitable research agendas in the AI field to attract unwitting researchers and practitioners, as Timnit Gebru and Émile P. Torres describe:
@@
Another line of historical lens traces the imperial, capitalist, and colonial roots in AI that questions the “origin myths that treat AI as a natural marriage of logic and computing in the mid-twentieth century” (Ali et al., 2023). Ali et al. (2023) in their editorial Histories of artificial intelligence: a genealogy of power illustrate that the purported different periods in AI history have “shared logics – especially managerial, military, industrial and computational – cut across them, often in ways that reinforce oppressive racial and gender hierarchies.” In the paper Animo nullius: on AI’s origin story and a data colonial doctrine of discovery (Penn, 2023), Jonnie Penn “traces elements of the theoretical origins of artificial intelligence to capitalism, not neurophysiology,” and compares data colonialism and privatization of the data commons with historical colonialism. The enclosure of the data commons is naturalized and veiled via the scientific pedigree of AI. In his book Artificial Whiteness: Politics and Ideology in Artificial Intelligence, Yarden Katz inspects the history of AI and its formation in the American military-industrial complex. The nebulous concept of AI helps its promoters to continuously reinvent and rebrand it to serve the aims of empire and capital (Katz, 2020). In the paper The Problem with Intelligence: Its Value-Laden History and the Future of AI, Stephen Cave summarizes the value-laden history of intelligence, and the entanglement of intelligence with the matrices of domination, patriarchy, colonialism, scientific racism, and eugenics (Cave, 2020).
@@
B.2 Section 2. The mainstream imagination of AI (MIA) is not as ethical as we assume
B.2.1 S1. AI can be omniscient
@@
In their book The Blind Spot: Why Science Cannot Ignore Human Experience (Frank et al., 2024), Adam Frank, Evan Thompson, and Marcelo Gleiser argue that the crises of modern science and technology lie in the underlying problematic philosophy and worldview, which they name the Blind Spot worldview. The Blind Spot worldview is a “culturally ubiquitous mind-set”, “like the air we breathe.”
‘[F]or most people, including most scientists, [the Blind Spot] it’s so pervasive that it doesn’t seem like philosophy at all. Rather, people think it’s just “what science says.”… In truth, however, it’s not what science says. Instead, it’s an optional metaphysics attached to, but separable from, the actual practice of science. There are other and better ways to understand the relationship between science and the world, as we argue here.’ [^15] From Frank et al. (2024) Chapter 1 The Surreptitious Substitution: Philosophical Origins of the Blind Spot ‘We call the source of the meaning crisis the Blind Spot. At the heart of science lies something we do not see that makes science possible, just as the blind spot lies at the heart of our visual field and makes seeing possible. In the visual blind spot sits the optic nerve; in the scientific blind spot sits direct experience—that by which anything appears, shows up, or becomes available to us. It is a precondition of observation, investigation, exploration, measurement, and justification. Things appear and become available thanks to our bodies and their feeling and perceiving capacities. Direct experience is bodily experience. “The body is the vehicle of being in the world,” says French philosopher Maurice Merleau-Ponty, but as we will see, firsthand bodily experience lies hidden in the Blind Spot.’ [^16] From Frank et al. (2024) Chapter An Introduction to the Blind Spot “We have argued that we must inscribe ourselves back into the scientific narrative as its creators. Science rests on how we experience the world. There is no way to take ourselves out of the story and tell it from a God’s-eye perspective. Forgetting this fact means succumbing to the Blind Spot, and that means losing our way in both science and all the critical ways that science shapes society.” [^17] From Frank et al. (2024) Chapter Afterword
In the later part of the Introduction of the book, Frank et al. (2024) provide a parable of the bodily experience of warm and cold and the scientific concept and measurement of temperature to illustrate the Blind Spot worldview:
@@
Frank et al. (2024) then return to the example of temperature to illustrate how these pathologies manifest in this example:
@@
In book Chapter 7 Cognition, Frank et al. (2024) illustrate how the Blind Spot worldview manifests in the AI field, which they name the computational blind spot:
@@
The MIA on AI’s omniscient prospect can be illustrated from the following excerpt in the book Artificial Whiteness: Politics and Ideology in Artificial Intelligence (Katz, 2020): ‘An industry of experts from the academic and policy worlds has emerged to interpret AI’s significance. According to the experts, AI could grow the economy and free people from labor burdens, help address climate change, remove the “bias” from the courts of law, automate scientific discovery, reinvent journalism (and hence democracy), and potentially produce, not long from now, “superhuman” intelligent machines that may launch new “civilizations” across the galaxy.’
@@
An example of ignoring human experience and denying human’s indispensable epistemic footing in AI is the myth of value-free and theory-free AI, which state that AI model and its data can be free from their contexts, human values, knowledge, theories, and interventions to represent the ideal neutral and objective knowledge, a “view from nowhere” (Nagel, 1989). For example, Chin-Yee & Upshur (2019) point out that ‘[p]roponents of big data and machine learning retort that the “atheoretical” perspective is in fact a strength of these methods, avoiding preconceptions that stifle scientific progress. Some pundits have gone as far as to declare the “end of theory” in the era of big data (Anderson 2008).’
@@
The contemporary development in epistemology, especially feminist epistemology, argues against this outmoded perspective. Discussions are usually under the term of “theory-ladenness of observation” (Chin-Yee & Upshur, 2019). In her influential paper Situated Knowledges: The Science Question in Feminism and the Privilege of Partial Perspective (Haraway, 1988), Donna Haraway denies the knowledge of “God trick” “of being nowhere while claiming to see comprehensively,” and instead proposes the feminist objectivity of situated knowledge, the knowledge from partial perspective:
@@
‘Above all, rational knowledge does not pretend to disengagement: to be from everywhere and so nowhere, to be free from interpretation, from being represented, to by fully self-contained or fully formalizable. Rational knowledge is a process of ongoing critical interpretation among “fields” of interpreters and decoders. Rational knowledge is power-sensitive conversation. … So science becomes the paradigmatic model, not of closure, but of that which is contestable and contested. Science becomes the myth, not of what escapes human agency and responsibility in a realm above the fray, but, rather, of accountability and responsibility for translations and solidarities linking the cacophonous visions and visionary voices that characterize the knowledges of the subjugated. … We seek not the knowledges ruled by phallogocentrism (nostalgia for the presence of the one true Word) and disembodied vision. We seek those ruled by partial sight and limited voice—not partiality for its own sake but, rather, for the sake of the connections and unexpected openings situated knowledges make possible. Situated knowledges are about communities, not about isolated individuals. The only way to find a larger vision is to be somewhere in particular. The science question in feminism is about objectivity as positioned rationality. Its images are not the products of escape and transcendence of limits (the view from above) but the joining of partial views and halting voices into a collective subject position that promises a vision of the means of ongoing finite embodiment, of living within limits and contradictions—of views from somewhere.’
@@
Because everyone has a unique perspective and contribution to collective knowledge, downplaying the role of some knowers in favor of other knowers is the manifestation of epistemic injustice (Fricker, 2007; Symons & Alvarado, 2022). For example, stating that AI can replace humans is to downplay the epistemic footing of human experience for AI in favor of the epistemic achievement of the AI models and their creators; Failure to acknowledge the intrinsic epistemic limitation of AI that cannot represent all possible perspectives and knowledges is to exclude the knowledges and experience of knowers (usually the marginalized groups) that are not included in AI modeling in favor of knowers whose knowledges are included in the AI model. The latter case is manifested in her book Artificial Knowing: Gender and the Thinking Machine (Adam, 1998), where sociologist Alison Adam argues that the “view from nowhere” in AI actually disguises the white, male, middle-class perspective as the natural, universal, and gold standard. Furthermore, Andrews et al. (2024) argue how this value- and theory-free myth has underpinned pseudoscientific and unethical practice in AI.
@@
Complex systems provide another perspective to understand why knowledge is situated. In the book Critical complexity (Cilliers, 2016), Paul Cilliers describes characteristics of complex systems that consist of a large number of elements that dynamically interact with each other. The interactions are rich, nonlinear, usually have a short range, and there are loops in the interactions. Complex systems are usually open systems, far from equilibrium, and have a history. Since human knowledge and AI modeling often aim to deal with complex systems, understanding complex systems can help us better understand knowledge, as described in (Cilliers, 2016):
@@
‘In order to predict the behaviour of a system accurately, we need a detailed understanding of that system, i.e., a model. Since the nature of a complex system is the result of the relationships distributed all over the system, such a model will have to reflect all these relationships. Since they are nonlinear, no set of interactions can be represented by a set smaller than the set itself – superposition does not hold. This is one way of saying that complexity is not compressible. Moreover, we cannot accurately determine the boundaries of the system, because it is open. In order to model a system precisely, we therefore have to model each and every interaction in the system, each and every interaction with the environment – which is of course also complex – as well as each and every interaction in the history of the system. In short, we will have to model life, the universe and everything. There is no practical way of doing this.’ [^18] From Cilliers (2016) Chapter What can we learn from a theory of complexity? ‘If one acknowledges the complexity of a system, it becomes more difficult to talk about “natural” boundaries. Boundaries are still required if we want to talk about complex systems in a meaningful way – they are in fact necessary, as argued above – but there are strategic considerations at stake when drawing them. These considerations may include subjective, or intersubjective components, but this does not mean that they are arbitrary. A complex system has structure and patterns that would render some descriptions more meaningful than others, but the point is that we do not have an a priori decision procedure for determining when we are dealing with something “more meaningful”. The contingent and historic nature of complex systems entails that our understanding of the system will have to be continually revised; the frames of our models will have to change. The boundaries of complex systems cannot be identified objectively, finally and completely. This supports the argument that our knowledge of complex systems cannot be reduced to formal algorithms, but has to incorporate considerations of what the knowledge is for. The criteria used to evaluate the knowledge are not independent things; they co-determine the nature of the knowledge (see Rosen 1996). Knowledge cannot be abstract and complete – we cannot “know” something like that. For us to have knowledge about something, it has to be limited.’ [^19] From Cilliers (2016) Chapter Knowledge, limits and boundaries
B.2.2 S2. Replacing human with AI can free human
@@
In addition to a series of labor displacements described in the main text, Acemoglu & Johnson (2024) describe in the early stages of the Industrial Revolution, with the introduction of power-loom factories, women and child labor were used to replace craftsmen, which had resulted in the same consequences of cutting wages and gaining more control over workers.
@@
A number of works have shown how the hidden human labor is disguised as the “magic”, “intelligence,” or “progress” of machines. Neda Atanasoski and Kalindi Vora have a precise description in their book Surrogate Humanity: Race, Robots, and the Politics of Technological Futures (Atanasoski & Vora, 2019b):
@@
Harry Law also reveals the concealment of labor in the history of early artificial neural networks built in Bell Labs, in the paper Bell Labs and the ‘neural’ network, 1986–1996 (Law, 2023):
@@
The main text discussion of the two groups of people who are replaced by AI and who work behind AI can fit into the big picture of the “institutionalized societal order” of capitalism (Fraser, 2022). In her book Cannibal Capitalism: How our System is Devouring Democracy, Care, and the Planet – and What We Can Do About It, philosopher Nancy Fraser provides a framework to deepen our understanding of the capitalist society. In the framework, she expands the foreground economic features of capitalism to their background non-economic preconditions, including “social reproduction, the earth’s ecology, political power, and ongoing infusions of wealth expropriated from racialized peoples”  (Fraser, 2022). She explores three perspectives of capitalism: exchange, exploitation, and expropriation. The market exchange perspective, characterized by self-interest transactors and growth-and-efficiency-maximizing logic, only narrowly defines capitalism by ignoring its structural dependence on the above background conditions. The backstory of the front story of exchange is exploitation, where workers’ surplus labor time is appropriated by capital. The backstory of exploitation is expropriation that makes exploitation possible and profitable, where capitalists brutally confiscate people’s assets, including labor, land, minerals, and energy resources. Exploitation and expropriation create two groups of workers, as Fraser states:
@@
Regarding how automation impoverishes work, Jason Resnikoff has extended evidence and arguments in his book Labor’s End: How the Promise of Automation Degraded Work (Resnikoff, 2021).
B.2.3 S3. AI can promote productivity and prosperity
Regarding Daron Acemoglu and Simon Johnson’s review of the economic history and technical change from the Middle Age to present, we summarize two examples in their book Power and Progress: Our Thousand-Year Struggle Over Technology and Prosperity to show how they associate the changes in workers’ wages, welfare, working conditions, and directions of technical development with changes in social and political power (Acemoglu & Johnson, 2023).
The first example is described in Chapter 6 Casualties of Progress. In it Acemoglu & Johnson (2023) describe that in Britain, the productivity gain from the early stages of the Industrial Revolution was not shared with workers with uncontrolled pollution and exacerbated public health, until later when workers struggled for political power such as union legalization and the right to vote.
@@
B.2.4 S4. Harms of AI are due to malicious use
Regarding the discourse that harms of AI are often due to unintended or malicious use, Nassim Parvin and Anne Pollock analyze the political implications of such a discourse in their paper Unintended by Design: On the Political Uses of “Unintended Consequences” (Parvin & Pollock, 2020):

### v2

@@ 1. The phenomenon of monoculture in the AI community, and why questioning the objectives of AI matters?
Historical studies show that the AI field did not develop in a vacuum, but was closely related to the social, economic, and political power, ideologies, and values at a given time. Scholars have traced the linkage between AI and ideologies such as eugenics, racism, colonialism, anti-democratic and anti-humanist utopianism [31, 12]. It should be noted that although the majority in the AI community may not be aware of or do not agree with these ideologies, it is the power and historical contingency that enabled a few privileged people, not the public or the majority in the community via a democratic process, to influence the research agenda and the underlying ideologies that set the paths of least resistance in the AI community. As the previous part indicates, once the rules of the game are set, we can know where the game is going without any information about the game players. And the rules of the game can appear to be necessary, obvious, natural, and inevitable. As Jenna Burrell and Jacob Metcalf put it in their editorial Introduction for the special issue of “Ideologies of AI and the consolidation of power”: Naming power [31]:
@@
A number of works have identify the goal of AI and its roots in eugenics. In their book How Data Happened: A History from the Age of Reason to the Age of Algorithms [107], Chris Wiggins and Matthew L. Jones trace the historical genesis of the contemporary statistics and data science in figures such as Francis Galton, who is the originator of eugenics, and Karl Pearson, who institutionalized eugenics. In her book Discriminating Data: Correlation, Neighborhoods, and the New Politics of Recognition [35], Wendy Hui Kyong Chun reveals the ties between correlation in machine learning and its eugenic history: both “seek to tie the past to the future—correlation to prediction—through supposedly eternal, unchanging biological attributes.” In their paper The TESCREAL bundle: Eugenics and the promise of utopia through artificial general intelligence [52], Timnit Gebru and Émile P. Torres trace the field’s current unquestioned motivation for artificial general intelligence (AGI) back to the normative framework of the newer version of eugenics pushed by a few powerful leading figures. The eugenic ideals are dubbed the “TESCREAL bundle,” which denotes “transhumanism, Extropianism, singularitarianism, (modern) cosmism, Rationalism, Effective Altruism, and longtermism.” These utopia goals justify the creation of unsafe algorithms, concentration of power, evasion of accountability, and ongoing harms to marginalized groups upon which AI is built. Furthermore, such ideologies are naturalized so that they appear as inevitable research agendas in the AI field to attract unwitting researchers and practitioners, as Timnit Gebru and Émile P. Torres describe:
@@
Another line of historical lens traces the imperial, capitalist, and colonial roots in AI that questions the “origin myths that treat AI as a natural marriage of logic and computing in the mid-twentieth century” [12]. Ali et al., [12] in their editorial Histories of artificial intelligence: a genealogy of power illustrate that the purported different periods in AI history have “shared logics – especially managerial, military, industrial and computational – cut across them, often in ways that reinforce oppressive racial and gender hierarchies.” In the paper Animo nullius: on AI’s origin story and a data colonial doctrine of discovery [89], Jonnie Penn “traces elements of the theoretical origins of artificial intelligence to capitalism, not neurophysiology,” and compares data colonialism and privatization of the data commons with historical colonialism. The enclosure of the data commons is naturalized and veiled via the scientific pedigree of AI. In his book Artificial Whiteness: Politics and Ideology in Artificial Intelligence, Yarden Katz inspects the history of AI and its formation in the American military-industrial complex. The nebulous concept of AI helps its promoters to continuously reinvent and rebrand it to serve the aims of empire and capital [65]. In the paper The Problem with Intelligence: Its Value-Laden History and the Future of AI, Stephen Cave summarizes the value-laden history of intelligence, and the entanglement of intelligence with the matrices of domination, patriarchy, colonialism, scientific racism, and eugenics [33].
@@
D.2 Section 2. The mainstream imagination of AI (MIA) is not as ethical as we assume
D.2.1 S1. AI can be capable of solving many hard problems
@@
In their book The Blind Spot: Why Science Cannot Ignore Human Experience [47], Adam Frank, Evan Thompson, and Marcelo Gleiser argue that the crises of modern science and technology lie in the underlying problematic philosophy and worldview, which they name the Blind Spot worldview. The Blind Spot worldview is a “culturally ubiquitous mind-set”, “like the air we breathe.”
‘[F]or most people, including most scientists, [the Blind Spot] it’s so pervasive that it doesn’t seem like philosophy at all. Rather, people think it’s just “what science says.”… In truth, however, it’s not what science says. Instead, it’s an optional metaphysics attached to, but separable from, the actual practice of science. There are other and better ways to understand the relationship between science and the world, as we argue here.’ [^17] From Frank et al., [47] Chapter 1 The Surreptitious Substitution: Philosophical Origins of the Blind Spot ‘We call the source of the meaning crisis the Blind Spot. At the heart of science lies something we do not see that makes science possible, just as the blind spot lies at the heart of our visual field and makes seeing possible. In the visual blind spot sits the optic nerve; in the scientific blind spot sits direct experience—that by which anything appears, shows up, or becomes available to us. It is a precondition of observation, investigation, exploration, measurement, and justification. Things appear and become available thanks to our bodies and their feeling and perceiving capacities. Direct experience is bodily experience. “The body is the vehicle of being in the world,” says French philosopher Maurice Merleau-Ponty, but as we will see, firsthand bodily experience lies hidden in the Blind Spot.’ [^18] From Frank et al., [47] Chapter An Introduction to the Blind Spot “We have argued that we must inscribe ourselves back into the scientific narrative as its creators. Science rests on how we experience the world. There is no way to take ourselves out of the story and tell it from a God’s-eye perspective. Forgetting this fact means succumbing to the Blind Spot, and that means losing our way in both science and all the critical ways that science shapes society.” [^19] From Frank et al., [47] Chapter Afterword
In the later part of the Introduction of the book, Frank et al., [47] provide a parable of the bodily experience of warm and cold and the scientific concept and measurement of temperature to illustrate the Blind Spot worldview:
@@
Frank et al., [47] then return to the example of temperature to illustrate how these pathologies manifest in this example:
@@
In book Chapter 7 Cognition, Frank et al., [47] illustrate how the Blind Spot worldview manifests in the AI field, which they name the computational blind spot:
@@
The MIA on AI’s omniscient prospect can be illustrated from the following excerpt in the book Artificial Whiteness: Politics and Ideology in Artificial Intelligence [65]: ‘An industry of experts from the academic and policy worlds has emerged to interpret AI’s significance. According to the experts, AI could grow the economy and free people from labor burdens, help address climate change, remove the “bias” from the courts of law, automate scientific discovery, reinvent journalism (and hence democracy), and potentially produce, not long from now, “superhuman” intelligent machines that may launch new “civilizations” across the galaxy.’
@@
An example of ignoring human experience and denying human’s indispensable epistemic footing in AI is the myth of value-free and theory-free AI, which state that AI model and its data can be free from their contexts, human values, knowledge, theories, and interventions to represent the ideal neutral and objective knowledge, a “view from nowhere” [81]. For example, Chin-Yee and Upshur, [34] point out that ‘[p]roponents of big data and machine learning retort that the “atheoretical” perspective is in fact a strength of these methods, avoiding preconceptions that stifle scientific progress. Some pundits have gone as far as to declare the “end of theory” in the era of big data (Anderson 2008).’
@@
The contemporary development in epistemology, especially feminist epistemology, argues against this outmoded perspective. Discussions are usually under the term of “theory-ladenness of observation” [34]. In her influential paper Situated Knowledges: The Science Question in Feminism and the Privilege of Partial Perspective [58], Donna Haraway denies the knowledge of “God trick” “of being nowhere while claiming to see comprehensively,” and instead proposes the feminist objectivity of situated knowledge, the knowledge from partial perspective:
@@
‘Above all, rational knowledge does not pretend to disengagement: to be from everywhere and so nowhere, to be free from interpretation, from being represented, to by fully self-contained or fully formalizable. Rational knowledge is a process of ongoing critical interpretation among “fields" of interpreters and decoders. Rational knowledge is power-sensitive conversation. … So science becomes the paradigmatic model, not of closure, but of that which is contestable and contested. Science becomes the myth, not of what escapes human agency and responsibility in a realm above the fray, but, rather, of accountability and responsibility for translations and solidarities linking the cacophonous visions and visionary voices that characterize the knowledges of the subjugated. … We seek not the knowledges ruled by phallogocentrism (nostalgia for the presence of the one true Word) and disembodied vision. We seek those ruled by partial sight and limited voice—not partiality for its own sake but, rather, for the sake of the connections and unexpected openings situated knowledges make possible. Situated knowledges are about communities, not about isolated individuals. The only way to find a larger vision is to be somewhere in particular. The science question in feminism is about objectivity as positioned rationality. Its images are not the products of escape and transcendence of limits (the view from above) but the joining of partial views and halting voices into a collective subject position that promises a vision of the means of ongoing finite embodiment, of living within limits and contradictions—of views from somewhere.’
@@
Because everyone has a unique perspective and contribution to collective knowledge, downplaying the role of some knowers in favor of other knowers is the manifestation of epistemic injustice [50, 101]. For example, stating that AI can replace humans is to downplay the epistemic footing of human experience for AI in favor of the epistemic achievement of the AI models and their creators; Failure to acknowledge the intrinsic epistemic limitation of AI that cannot represent all possible perspectives and knowledges is to exclude the knowledges and experience of knowers (usually the marginalized groups) that are not included in AI modeling in favor of knowers whose knowledges are included in the AI model. The latter case is manifested in her book Artificial Knowing: Gender and the Thinking Machine [9], where sociologist Alison Adam argues that the “view from nowhere” in AI actually disguises the white, male, middle-class perspective as the natural, universal, and gold standard. Furthermore, Andrews et al., [13] argue how this value- and theory-free myth has underpinned pseudoscientific and unethical practice in AI.
@@
Complex systems provide another perspective to understand why knowledge is situated. In the book Critical complexity [36], Paul Cilliers describes characteristics of complex systems that consist of a large number of elements that dynamically interact with each other. The interactions are rich, nonlinear, usually have a short range, and there are loops in the interactions. Complex systems are usually open systems, far from equilibrium, and have a history. Since human knowledge and AI modeling often aim to deal with complex systems, understanding complex systems can help us better understand knowledge, as described in [36]:
@@
‘In order to predict the behaviour of a system accurately, we need a detailed understanding of that system, i.e., a model. Since the nature of a complex system is the result of the relationships distributed all over the system, such a model will have to reflect all these relationships. Since they are nonlinear, no set of interactions can be represented by a set smaller than the set itself – superposition does not hold. This is one way of saying that complexity is not compressible. Moreover, we cannot accurately determine the boundaries of the system, because it is open. In order to model a system precisely, we therefore have to model each and every interaction in the system, each and every interaction with the environment – which is of course also complex – as well as each and every interaction in the history of the system. In short, we will have to model life, the universe and everything. There is no practical way of doing this.’ [^20] From Cilliers, [36] Chapter What can we learn from a theory of complexity? ‘If one acknowledges the complexity of a system, it becomes more difficult to talk about “natural” boundaries. Boundaries are still required if we want to talk about complex systems in a meaningful way – they are in fact necessary, as argued above – but there are strategic considerations at stake when drawing them. These considerations may include subjective, or intersubjective components, but this does not mean that they are arbitrary. A complex system has structure and patterns that would render some descriptions more meaningful than others, but the point is that we do not have an a priori decision procedure for determining when we are dealing with something “more meaningful”. The contingent and historic nature of complex systems entails that our understanding of the system will have to be continually revised; the frames of our models will have to change. The boundaries of complex systems cannot be identified objectively, finally and completely. This supports the argument that our knowledge of complex systems cannot be reduced to formal algorithms, but has to incorporate considerations of what the knowledge is for. The criteria used to evaluate the knowledge are not independent things; they co-determine the nature of the knowledge (see Rosen 1996). Knowledge cannot be abstract and complete – we cannot “know” something like that. For us to have knowledge about something, it has to be limited.’ [^21] From Cilliers, [36] Chapter Knowledge, limits and boundaries
D.2.2 S2. Replacing human with AI can free human
@@
In addition to a series of labor displacements described in the main text, Acemoglu and Johnson, [3] describe in the early stages of the Industrial Revolution, with the introduction of power-loom factories, women and child labor were used to replace craftsmen, which had resulted in the same consequences of cutting wages and gaining more control over workers.
@@
A number of works have shown how the hidden human labor is disguised as the “magic”, “intelligence,” or “progress” of machines. Neda Atanasoski and Kalindi Vora have a precise description in their book Surrogate Humanity: Race, Robots, and the Politics of Technological Futures [15]:
@@
Harry Law also reveals the concealment of labor in the history of early artificial neural networks built in Bell Labs, in the paper Bell Labs and the ‘neural’ network, 1986–1996 [68]:
@@
The main text discussion of the two groups of people who are replaced by AI and who work behind AI can fit into the big picture of the “institutionalized societal order” of capitalism [48]. In her book Cannibal Capitalism: How our System is Devouring Democracy, Care, and the Planet – and What We Can Do About It, philosopher Nancy Fraser provides a framework to deepen our understanding of the capitalist society. In the framework, she expands the foreground economic features of capitalism to their background non-economic preconditions, including “social reproduction, the earth’s ecology, political power, and ongoing infusions of wealth expropriated from racialized peoples”  [48]. She explores three perspectives of capitalism: exchange, exploitation, and expropriation. The market exchange perspective, characterized by self-interest transactors and growth-and-efficiency-maximizing logic, only narrowly defines capitalism by ignoring its structural dependence on the above background conditions. The backstory of the front story of exchange is exploitation, where workers’ surplus labor time is appropriated by capital. The backstory of exploitation is expropriation that makes exploitation possible and profitable, where capitalists brutally confiscate people’s assets, including labor, land, minerals, and energy resources. Exploitation and expropriation create two groups of workers, as Fraser states:
@@
Regarding how automation impoverishes work, Jason Resnikoff has extended evidence and arguments in his book Labor’s End: How the Promise of Automation Degraded Work [93].
D.2.3 S3. AI can promote productivity and prosperity
Regarding Daron Acemoglu and Simon Johnson’s review of the economic history and technical change from the Middle Age to present, we summarize two examples in their book Power and Progress: Our Thousand-Year Struggle Over Technology and Prosperity to show how they associate the changes in workers’ wages, welfare, working conditions, and directions of technical development with changes in social and political power [2].
The first example is described in Chapter 6 Casualties of Progress. In it Acemoglu and Johnson, [2] describe that in Britain, the productivity gain from the early stages of the Industrial Revolution was not shared with workers with uncontrolled pollution and exacerbated public health, until later when workers struggled for political power such as union legalization and the right to vote.
@@
D.2.4 S4. Harms of AI are due to malicious use
Regarding the discourse that harms of AI are often due to unintended or malicious use, Nassim Parvin and Anne Pollock analyze the political implications of such a discourse in their paper Unintended by Design: On the Political Uses of “Unintended Consequences” [88]: